OpenStudy (voltage):

Look at the following numbers: −2, −1, 0, 1 Which pair of numbers has a sum of zero? 1, −1 −2, 0 −2, 1 −1, 0

OpenStudy (anonymous):

What happens when you add a negative with a positive?

OpenStudy (voltage):

idk

OpenStudy (anonymous):

When you add a negative and a positive together you will get a negative sum.

OpenStudy (anonymous):

the answer to this question is a simple as working out each problem.

OpenStudy (voltage):

oh

OpenStudy (anonymous):

So for the first choise you would do,\[1+ -1= x\]

OpenStudy (anonymous):

What would your answer be?

OpenStudy (mikeymaximum):

I got A.

OpenStudy (anonymous):

a postive plus a negative does not allways equal a negative a positive times a negative will always equal a negative though

OpenStudy (voltage):

ok

OpenStudy (anonymous):

But a is your answer because 1 +(-1)= 1-1= 0

OpenStudy (voltage):

thank you

OpenStudy (anonymous):

yep no problem